# Good involutions of conjugation subquandles

### v4

@@ Abstract.
As an application of our results, we implement group-theoretic algorithms that compute all good involutions of conjugation quandles and core quandles; we provide data for those up to order $23$. As another application, we construct infinite families of connected, noninvolutive symmetric quandles.
We also classify symmetric and Legendrian racks, quandles, and kei up to order $8$ using a computer search. Finally, we exhibit an equivalence of categories between racks and Legendrian racks that induces an equivalence between involutory racks, Legendrian kei, and symmetric kei.
@@ 1.1. Overview of results
In particular, if $\operatorname{Conj}X$ is not self-dual or if there exists some $x\in X$ such that $xy\notin Z(\langle X\rangle)$ for all $y\in X$, then $\operatorname{Conj}X$ has no good involutions. (See Corollary 4.4 and Proposition 6.4.)
@@
Let $G$ be a group such that the core quandle $\operatorname{Core}G=(G,s)$ is nontrivial, and let $A$ be the $2$-torsion subgroup of the center $Z(G)$. Then for all functions $\rho:G\to G$, $\rho$ is a good involution of $\operatorname{Core}G$ if and only if there exists a function $\alpha:G\to A$ such that
@@
for all $g\in G$. (See Theorem 7.5; cf. Remark 7.6 and Corollaries 7.9 and 7.10.)
@@
As applications of Theorem 1.3 and 1.4, we describe procedures (Algorithms 8.1 and 8.5) for constructing all good involutions of conjugation subquandles and core quandles, and we provide data for groups up to order $23$ in a GitHub repository [code]. This addresses Problem 1.2.
@@ 1.1.1. Computer search
Table 1. Enumeration of isomorphism classes of various types of symmetric racks up to order $8$.
@@
Through an exhaustive search in the computer algebra system GAP [GAP4], we tabulated all isomorphism classes of symmetric racks, quandles, and kei up to order $9$. This answers Problem 1.2 for all racks up to order $9$. Our classification used Vojtěchovský and Yang’s [library] library of racks in a way similar to the algorithms described in [taGL]*App. A.
We provide our code and data in a GitHub repository [code]. Table 1 enumerates our results.
@@ 1.1.2. CNS-quandles
Connected, noninvolutory symmetric quandles, which we call CNS-quandles, are particularly helpful for studying $3$-manifolds [nosaka] and nonorientable surface-knots [dihedral]. However, CNS-quandles are difficult to construct. For example, generalized Alexander quandles have good involutions if and only if they are kei [taniguchi]*Thm. 1.3, and our computer search showed the following.
@@
Up to isomorphism, the only CNS-quandle of order $8$ or lower is the symmetric Galkin quandle of order $6$ constructed by Clark et al. [galkin-symm]*Prop. 5.13.
Thus, the following problems arise.
@@
Carter et al. [dihedral]*Thm. 3.1 constructed CNS-quandles of order $2^{m}m$ for all odd integers $m\geq 1$, while Clark et al. [galkin-symm]*Prop. 5.13 constructed CNS-quandles of order $6n$ for all positive integers $n\geq 1$. To the author’s knowledge, these remain the only two answers to Problems 1.6 and 1.7 in the literature. To address this, we introduce a broad family of CNS-quandles in Proposition 9.1. In particular, two special cases of our construction give answers to Problem 1.7.
@@
If $q=3=k$, then for all integers $m\geq 3$ and $n\geq 5$ such that $n\not\equiv 0,1,2$ mod $9$, the CNS-quandles in Proposition 1.8 do not share orders with the CNS-quandles constructed by Clark et al. Most of them also do not share orders with the CNS-quandles constructed by Carter et al.
@@ 1.1.3. Connections with Legendrian racks
Generalizing prior work of Kulkarni and Prathamesh [original] in 2017, Ceniceros et al. [ceniceros] introduced Legendrian racks in 2021. In 2023, Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced generalizations of Legendrian racks called GL-racks; in 2025, the author [taGL] studied algebraic aspects of GL-racks and classified them up to order $8$.
@@ 1.2. Structure of the paper
In Sections 2 and 3, we provide definitions and important examples of racks, quandles, kei, rack (anti)automorphisms, and good involutions.
@@
In Section 5, we address Problems 1.1 and 1.2 for faithful racks. In particular, we show that Galkin quandles have no good involutions other than those constructed by Clark et al. [galkin-symm]*Prop. 5.13. This shows that infinitely many self-dual quandles have no good involutions.
@@
In Section 8, we apply Theorems 1.3 and 1.4 to describe procedures for computing all good involutions of conjugation subquandles and core quandles. Using an implementation of these procedures in GAP [GAP4], we provide data for groups up to order $23$. This addresses Problem 1.2.
In Section 9, we consider several examples of connected conjugation subquandles and compute all of their good involutions. Along the way, we construct infinite families of CNS-quandles; in particular, we obtain Proposition 1.8. This addresses Problems 1.6 and 1.7.
@@ 2.1. Racks and quandles
Note that $\operatorname{Core}G$ is faithful if and only if the center $Z(G)$ is $2$-torsionless [core]*Lem. 4.1. Moreover, $\operatorname{Core}G$ is a permutation rack if and only if $G$ is elementary $2$-abelian, in which case $\operatorname{Core}G$ is a trivial quandle.
Furthermore, if $G$ is an additive abelian group, then the core quandle $T(G):=\operatorname{Core}(G)$ is called a Takasaki kei [quandlebook]*Ex. 54. Note that $T(G)$ is faithful if and only if $G$ is $2$-torsionless.
@@
Let $G$ be a group, and let $\varphi\in\operatorname{Aut}G$ be an automorphism of $G$. The generalized Alexander quandle $\operatorname{GAlex}(G,\varphi)$ is defined to be the quandle $(G,s)$, where
@@ 2.1.1. Conjugation subquandles
We discuss subquandles of conjugation quandles. In the following, let $G$ be a group, and let $s:G\to S_{G}$ be the rack structure of $\operatorname{Conj}G$.
@@ 2.1.2. Rack homomorphisms
Given two racks $(X,s)$ and $(Y,t)$, we say that a map $\varphi:X\to Y$ is a rack homomorphism if
@@
If $\operatorname{Conj}X$ is a subquandle of $\operatorname{Conj}G$, then $\operatorname{Conj}X$ is connected if and only if $X$ is a conjugacy class in the subgroup $H:=\langle X\rangle\leq G$.
@@ 3.1. Rack antiautomorphisms
An antiautomorphism of $R$ is a rack isomorphism $\iota:R\xrightarrow{\sim}R^{\mathrm{op}}$. Let $\operatorname{Anti}R$ be the set of antiautomorphisms of $R$. We say that $R$ is self-dual if $\operatorname{Anti}R$ is nonempty.
@@ 3.2. Good involutions
Let $G$ be a group, and let $z\in Z(G)$ be a central element of order $1$ or $2$. Then multiplication by $z$ is a good involution of $\operatorname{Core}G$. See Theorem 7.5 for a generalization of this example.
@@ 4. General results for racks
Let $R=(X,s)$ be a rack. In this section, we characterize good involutions $\rho\in\operatorname{Good}R$ as involutive central antiautomorphisms of $R$ (Proposition 4.2). As a result, self-duality is a necessary condition for $R$ to have good involutions (Corollary 4.4), and every good involution factors through some automorphism $\varphi$ and a given antiautomorphism $\iota$ (Corollary 4.7).
@@ 4.2. Factoring good involutions
Since $R\cong R^{\mathrm{op}}$, this follows from Observation 3.2 and the definition of $\operatorname{Anti}R$. ∎
@@
for all odd $3\leq k<n$. Then $R:=(X,s)$ is a kei, so Example 3.3 implies that $\operatorname{Anti}R$ is the elementary abelian $2$-group
@@ 5.1. General results
Let $\rho_{1},\rho_{2}\in\operatorname{Good}R$. For all $x\in X$,
$s_{\rho_{1}(x)}=s^{-1}_{x}=s_{\rho_{2}(x)},$
so $\rho_{1}(x)=\rho_{2}(x)$ because $R$ is faithful. Therefore, $\rho_{1}=\rho_{2}$. ∎
@@
An inverse to Proposition 5.3 holds; see Corollary 6.5.
Proposition 5.5.
@@
Let $G$ be a group, and let $\varphi\in\operatorname{Aut}G$ be an automorphism of $G$. By a theorem of Taniguchi [taniguchi]*Thm. 1.3, the generalized Alexander quandle $\operatorname{GAlex}(G,\varphi)$ has a good involution if and only if it is a kei. In this case, if the only fixed point of $\varphi$ is the identity element $1\in G$, then Example 2.8 and Proposition 5.5 imply that $\operatorname{Good}(\operatorname{GAlex}(G,\varphi))=\{\operatorname{id}_{G}\}$.
In particular, if $G$ is abelian and $2$-torsionless, then the only good involution of the Takasaki kei $T(G)$ is the identity map $\operatorname{id}_{G}$; cf. Corollary 7.10. This generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles $R_{n}$ of odd order.
@@ 5.2. Symmetric Galkin quandles
In particular, Clark et al. [galkin-symm]*Prop. 5.13 constructed good involutions of certain Galkin quandles. We show that no other good involutions of Galkin quandles exist. Aside from addressing Problems 1.1 and 1.2, this is also motivated by finding infinitely many quandles that have involutive antiautomorphisms and no good involutions; cf. Remark 4.5.
@@
$\mu(n):=\begin{cases}2&\text{if }n=0,\\ -1&\text{if }n=1,2,\end{cases}\qquad\tau(n):=\begin{cases}0&\text{if }n=0,1,\\ c,&\text{if }n=2.\end{cases}$
@@ 5.2.1. Results
Clark et al. [galkin-symm]*Prop. 5.13, cf. Cor. 5.12 constructed good involutions of certain Galkin quandles. Namely, let $c\in A$ be an element such that $02c=$ 0, and define
@@
If $02c=$ 0, then $\rho_{c}\in\operatorname{Good}(\mathcal{G}(A,c))$ by the construction of Clark et al. Conversely, for all good involutions $\rho=(\rho_{1},\rho_{2})$ of $\mathcal{G}(A,c)$ and for all $(x,a),(y,b)\in\mathbb{Z}/3\mathbb{Z}\times A$, the expression
@@
Taking $z$ to be $1$ and $2$ respectively yields
@@
from which we deduce that $02c=$ 0 and $\rho_{2}(b)=b+c$ for all $b\in A$. In particular, $\rho=\rho_{c}$. ∎
@@
If $A$ is $2$-torsionless, then for all $c\in A$, there are no good involutions of $\mathcal{G}(A,c)$.
@@ 6.2. Intermediate results
This follows from Propositions 5.3 and 6.4. ∎
@@
Consider the quandle $\operatorname{Conj}X$ from Example 2.16. Although $\operatorname{Conj}X$ is self-dual, Corollary 6.5 states that $\operatorname{Conj}X$ has no good involutions. This can also be seen from the fact that $\operatorname{Conj}X$ is a noninvolutory quandle of order $4$; Table 1 shows that no such quandles have good involutions.
@@ 6.3.1. Corollaries
yields the main claim due to Theorem 6.7. The final claim follows from Observation 6.10. ∎
@@
The bound $|{\operatorname{Good}(\operatorname{Conj}X)}|\leq|{\operatorname{Aut}(\operatorname{Conj}X)}|$ comes from Corollary 4.8. On the other hand, Remark 6.8 shows that there are at least as many functions $\zeta^{*}:X/H\to Z(H)$ as there are good involutions $\rho$; that is, $|{Z(H)}|^{k(X)}\geq|{\operatorname{Good}(\operatorname{Conj}X)}|$. The final claim follows from Observation 6.10. ∎
@@
By Corollary 6.5, the bounds in Corollaries 6.11 and 6.12 are equalities whenever $H$ is centerless and $\operatorname{Conj}X$ is inversion-closed. Furthermore, Example 9.2 provides a connected conjugation subquandle $\operatorname{Conj}X$ of order $30$ that attains these bounds even though $Z(H)$ is nontrivial, and Example 9.5 provides a similar example whose order is infinite.
@@ 7. Symmetric core quandles
In this section, we prove Theorem 1.4 (Theorem 7.5) and discuss three corollaries. This addresses Problems 1.1 and 1.2 for nontrivial core quandles $\operatorname{Core}G$.
Given a group $G$, recall that $\operatorname{Core}G$ is trivial if and only if $G$ is an elementary abelian $2$-group. Since good involutions of trivial quandles are already classified (see Example 3.10), we will assume throughout this section that $G$ is not an elementary abelian $2$-group. Let $A$ be the $2$-torsion subgroup of the center $Z(G)$:
$A:=\{z\in Z(G)\mid z^{2}=1\}.$

### v5

@@ Abstract.
As an application of our results, we implement group-theoretic algorithms that compute all good involutions of conjugation quandles and core quandles; we provide data for those up to order 232323. As another application, we construct infinite families of connected, noninvolutory symmetric quandles.
We also classify symmetric and Legendrian racks, quandles, and kei up to order 888 using a computer search. Finally, we exhibit an equivalence of categories between racks and Legendrian racks that induces an equivalence between involutory racks, Legendrian kei, and symmetric kei.
@@ 1. Introduction
This paper addresses Problems 1.1 and 1.2 for large classes of quandles of broad interest called conjugation quandles and core quandles. These quandles are constructed using the functors $\operatorname{Conj}$ and $\operatorname{Core}$ from the category of groups to the category of quandles. Joyce [joyce] and Matveev [matveev] independently introduced these quandles to generalize Wirtinger presentations of knot groups and develop complete invariants of knots.
@@ 1.1. Overview of results
Let $G$ be a group such that the core quandle $\operatorname{Core}G=(G,s)$ is nontrivial, and let $T$ be the subgroup
$T:=\{z\in Z(G)\mid z^{2}=1\}$
of the center $Z(G)$. Then for all functions $\rho:G\to G$, $\rho$ is a good involution of $\operatorname{Core}G$ if and only if there exists a function $\alpha:G\to T$ such that
@@
for all $g\in G$. (See Theorem 7.11; cf. Remark 7.13 and Corollaries 7.16 and 7.17.)
@@
As applications of Theorem 1.3 and 1.4, we describe procedures (Algorithms 8.1 and 8.5) for constructing all good involutions of conjugation subquandles and core quandles, and we provide data for groups up to order 232323 in a GitHub repository [code]. This addresses Problem 1.2.
@@ 1.1.1. Computer search
Table 1. Enumeration of isomorphism classes of various types of symmetric racks up to order 888.
@@
Through an exhaustive search in the computer algebra system GAP [GAP4], we tabulated all isomorphism classes of symmetric racks, quandles, and kei up to order 888. This answers Problem 1.2 for all racks up to order 888. Our classification used Vojtěchovský and Yang’s [library] library of racks in a way similar to the algorithms described in [taGL]*App. A.
We provide our code and data in a GitHub repository [code]. Table 1 enumerates our data.
@@ 1.1.2. CNS-quandles
Connected, noninvolutory symmetric quandles, which we call CNS-quandles, are particularly helpful for studying 333-manifolds [nosaka] and nonorientable surface-knots [dihedral]. However, CNS-quandles are difficult to construct. For example, generalized Alexander quandles have good involutions if and only if they are kei [taniguchi]*Thm. 1.3, and our computer search showed the following.
@@
Up to isomorphism, the only CNS-quandle of order 888 or lower is the symmetric Galkin quandle of order 666 constructed by Clark et al. [galkin-symm]*Prop. 5.13.
Thus, the following problems arise; see [galkin-symm]*Sec. 5.
@@
[…]
@@
If $q=3=k$, then for all integers $m\geq 3$ and $n\geq 5$ such that $n\not\equiv 0,1,2$ mod 999, the CNS-quandles in Proposition 1.8 do not share orders with the CNS-quandles constructed by Clark et al. Most of them also do not share orders with the CNS-quandles constructed by Carter et al.
@@ 1.1.3. Connections with Legendrian racks
Generalizing prior work of Kulkarni and Prathamesh [original] in 2017, Ceniceros et al. [ceniceros] introduced Legendrian racks in 2021. In 2023, Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced generalizations of Legendrian racks called GL-racks. In 2025, the author [taGL] studied algebraic aspects of GL-racks and classified them up to order 888, and Cheng and He [he] showed that GL-racks distinguish Legendrian knots up to the absolute values of their classical invariants.
@@ 1.2. Structure of the paper
In Section 2, we provide definitions and examples of racks, quandles, and kei. In Section 3, we do the same for rack antiautomorphisms and good involutions.
@@
In Section 5, we discuss Problems 1.1 and 1.2 for faithful racks. A class of examples shows that infinitely many quandles have involutory antiautomorphisms but no good involutions.
@@
In Section 8, we describe procedures for computing all good involutions of conjugation subquandles and core quandles. We provide data for groups up to order 232323 [code]. This addresses Problem 1.2.
In Section 9, we compute all good involutions of several classes of connected quandles. Along the way, we obtain Proposition 1.8. This addresses Problems 1.6 and 1.7.
@@ 2.1. Racks and quandles
Note that $\operatorname{Core}G$ is faithful if and only if the center $Z(G)$ is 222-torsionless [core2]*Lem. 4.1. Moreover, $\operatorname{Core}G$ is a permutation rack if and only if $G$ is an elementary abelian 222-group, in which case $\operatorname{Core}G$ is a trivial quandle.
Furthermore, if $G$ is an additive abelian group, then the core quandle $T(G):=\operatorname{Core}(G)$ is called a Takasaki kei [quandlebook]*Ex. 54. Note that $T(G)$ is faithful if and only if $G$ is 222-torsionless.
@@
[…]
@@ 2.1.1. Conjugation subquandles
We discuss subquandles of conjugation quandles. In the following, let $G$ be a group, and let $s:G\to S_{G}$ denote the rack structure of $\operatorname{Conj}G$.
@@ 2.1.2. Rack homomorphisms
[…]
@@
If $\operatorname{Conj}X$ is a subquandle of $\operatorname{Conj}G$, then $\operatorname{Conj}X$ is connected if and only if $X$ is a conjugacy class in the subgroup $\langle X\rangle\leq G$.
@@ 3.1. Rack antiautomorphisms
Definition 3.2.
Given a rack $R$, an antiautomorphism of $R$ is a rack isomorphism $\rho:R\xrightarrow{\sim}R^{\mathrm{op}}$. Let $\operatorname{Anti}R$ be the set of antiautomorphisms of $R$. We say that $R$ is self-dual if $\operatorname{Anti}R$ is nonempty.
@@ 3.2. Good involutions
Let $G$ be a group, and let $z\in Z(G)$ be a central element of order 111 or 222. Then multiplication by $z$ is a good involution of $\operatorname{Core}G$.
@@ 4. General results for racks
Let $R=(X,s)$ be a rack. In this section, we characterize good involutions $\rho\in\operatorname{Good}R$ as involutory central antiautomorphisms of $R$ (Proposition 4.2). As a result, self-duality is a necessary condition for $R$ to have good involutions (Corollary 4.4), and every good involution factors through some automorphism $\varphi$ and a given antiautomorphism $\iota$ (Corollary 4.7).
@@ 4.2. Factoring good involutions
Since $R\cong R^{\mathrm{op}}$, this follows from Observation 3.3 and the definition of $\operatorname{Anti}R$. ∎
@@
for all odd $3\leq k<n$. Then $R:=(X,s)$ is a kei, so Example 3.4 implies that $\operatorname{Anti}R$ is the elementary abelian 222-group
@@ 5.1. General results
Let $\rho,\rho^{\prime}\in\operatorname{Good}R$. For all $x\in X$,
$s_{\rho(x)}=s^{-1}_{x}=s_{\rho^{\prime}(x)},$
so $\rho(x)=\rho^{\prime}(x)$ because $R$ is faithful. Therefore, $\rho=\rho^{\prime}$. ∎
@@
An inverse to Corollary 5.3 holds; see Corollary 6.5.
Corollary 5.5.
@@
Let $G$ be a group, and let $\varphi\in\operatorname{Aut}G$ be an automorphism of $G$. By a theorem of Taniguchi [taniguchi]*Thm. 1.3, the generalized Alexander quandle $\operatorname{GAlex}(G,\varphi)$ has a good involution if and only if it is a kei. In this case, if the only fixed point of $\varphi$ is the identity element $1\in G$, then Example 2.8 and Corollary 5.5 imply that $\operatorname{Good}(\operatorname{GAlex}(G,\varphi))=\{\operatorname{id}_{G}\}$.
In particular, if $G$ is abelian and 222-torsionless, then the only good involution of the Takasaki kei $T(G)$ is the identity map $\operatorname{id}_{G}$; cf. Corollary 7.17. This generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles $R_{n}$ of odd order.
@@ 5.2. Symmetric Galkin quandles
In particular, Clark et al. [galkin-symm]*Prop. 5.13 constructed good involutions of certain Galkin quandles. We show that no other good involutions of Galkin quandles exist. Aside from addressing Problems 1.1 and 1.2, this is also motivated by finding infinitely many quandles that have involutory antiautomorphisms and no good involutions; cf. Remark 4.5.
@@
$\mu(n):=\begin{cases}2&\text{if }n=0,\\ -1&\text{if }n=1,2,\end{cases}\qquad\tau(n):=\begin{cases}0&\text{if }n=0,1,\\ c&\text{if }n=2.\end{cases}$
@@ 5.2.1. Results
[…]
@@
If $2c=0$, then $\rho_{c}\in\operatorname{Good}(\mathcal{G}(A,c))$ by the construction of Clark et al. Conversely, for all good involutions $\rho=(\rho_{1},\rho_{2})$ of $\mathcal{G}(A,c)$ and for all $(x,a),(y,b)\in\mathbb{Z}/3\mathbb{Z}\times A$, the expression
@@
Taking $z$ to be 111 and 222 respectively yields
@@
from which we deduce that $2c=0$ and $\rho_{2}(b)=b+c$ for all $b\in A$. In particular, $\rho=\rho_{c}$. ∎
@@
If $A$ is 222-torsionless, then for all $c\in A$, there are no good involutions of $\mathcal{G}(A,c)$.
@@ 6.2. Intermediate results
This follows from Corollary 5.3 and Proposition 6.4. ∎
@@
Consider the quandle $\operatorname{Conj}X$ from Example 2.16. Although $\operatorname{Conj}X$ is self-dual, Corollary 6.5 states that $\operatorname{Conj}X$ has no good involutions. This can also be seen from the fact that $\operatorname{Conj}X$ is a noninvolutory quandle of order 444; Table 1 shows that no such quandles have good involutions.
@@ 6.3.1. Corollaries
yields the main claim due to Theorem 6.7. The final claim follows from Example 6.8. ∎
@@
The bound $|{\operatorname{Good}(\operatorname{Conj}X)}|\leq|{\operatorname{Aut}(\operatorname{Conj}X)}|$ comes from Corollary 4.8. On the other hand, Remark 6.9 shows that there are at least as many functions $\zeta^{*}:X/H\to Z(H)$ as there are good involutions $\rho$; that is, $|{Z(H)}|^{k(X)}\geq|{\operatorname{Good}(\operatorname{Conj}X)}|$. The final claim follows from Example 6.8. ∎
@@
By Corollary 6.5, the bounds in Corollaries 6.11 and 6.12 are equalities whenever $H$ is centerless and $\operatorname{Conj}X$ is inversion-closed. Furthermore, Example 9.2 provides a connected conjugation subquandle $\operatorname{Conj}X$ of order 303030 that attains these bounds even though $Z(H)$ is nontrivial, and Example 9.5 provides a similar example whose order is infinite.
@@ 7. Symmetric core quandles
In this section, we prove Theorem 1.4 (Theorem 7.11) and discuss several related results. This addresses Problems 1.1 and 1.2 for nontrivial core quandles $\operatorname{Core}G$.
Given a group $G$, recall that $\operatorname{Core}G$ is trivial if and only if $G$ is an elementary abelian 222-group. Since good involutions of trivial quandles are already classified (see Example 3.11), we will let $G$ be a group of exponent greater than 222 throughout this section.